During the year ended December 31, 2020, Gluco Inc. split its stock on a 4-for-1 basis. In its annual report for 2019, the firm reported net income of $915,000 for 2019, with an average 258,000 shares of common stock outstanding for that year. There was no preferred stock.

  1. What amount of net income for 2019 will be reported in Gluco's 2020 annual report?
  2. Calculate Gluco's earnings per share for 2019 that would have been reported in the 2019 annual report. (Round your answer to 2 decimal places.)
  3. Calculate Gluco's earnings per share for 2019 that will be reported in the 2020 annual report for comparative purposes. (Round your answer to 2 decimal places
Answer #1
a Net Income for 2019 =$915,000
b EPS for 2019 =$915,000 / 258,000 shares =$3.55 per share
c Shares outstanding after stock-split =258,000*4 =1,032,000 shares
EPS =$915,000 / 1,032,000 =$0.89 per share